OpenStudy (agl202):

Check My answer? MEDAL!!! Identify all of the following solutions of square root of x plus 8 end root minus 6 equals x . x = −4 x = −7 x = −4 and x = −7 None of the above <----- My answer!

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[\sqrt{x+8}-6=x\]?

OpenStudy (misty1212):

there is one answer

OpenStudy (agl202):

\[\sqrt{x=2}\]

OpenStudy (agl202):

@misty1212

OpenStudy (misty1212):

\[x=-4\] works, you can check it by substitution q